# CEHR-GPT: A Scalable Multi-Task Foundation Model for Electronic Health Records

### v1

@@ 4.4 Generalization on EHR-SHOT Benchmark
We demonstrated generalizability of CEHR-GPT by evaluating it on the external ehrshot benchmarking dataset, which enables comparison with several competitive models available on the ehrshot Leaderboard, including MOTOR, Context, CLMBR, GBM, LR, RF, and GPT (Table 6 and Supplementary table 12). Overall, CEHR-GPT outperformed GPT in both fine-tuned and linear-probing settings. For Patient Outcomes, fine-tuning CEHR-GPT improved performance, with average gains of 1.2% in AUROC and 5.0% in AUPRC compared to linear probing. In contrast, on New Diagnosis tasks, fine-tuning CEHR-GPT yielded a higher AUROC with an average gain of 1.9% but slightly reduced AUPRC by 0.6%. Interestingly, GPT followed a different pattern: for Patient Outcomes, fine-tuning and linear probing performed similarly, with fine-tuning achieving a modest average improvement of 0.9% in AUPRC. However, for New Diagnosis, fine-tuning degraded performance relative to linear probing, particularly for two low-prevalence cohorts, Celiac and Pancreatic Cancer. A similar trend was observed for CEHR-GPT on Pancreatic Cancer, where linear probing outperformed fine-tuning. These results suggest that under extreme class imbalance, linear probing may offer a more robust approach for predicting rare events than fine-tuning.
More importantly, both fine-tuned and linear-probing versions of CEHR-GPT demonstrated strong performance across these tasks, achieving the best results on New Diagnosis prediction compared to all other models. Notably, CEHR-GPT outperformed models such as CLMBR, despite the latter being trained on data drawn from the same distribution (Stanford EHR data) as the ehrshot dataset. This suggests that simply extending the context window or adopting alternative architectures does not necessarily yield better performance on these tasks, and that CEHR-GPT’s novel patient representation design and time-specific learning objectives may be key contributors to its improved performance.
@@
Table 6: Performance metrics of selected models for patient outcome and new diagnosis predictions using the ehrshot dataset. This table summarizes both individual and average model performance across tasks. The Best Baseline column shows the best metric among CLMBR, GBM, LR, and RF. For MOTOR, only aggregated performance data is available since specific task details were not published on the ehrshot leaderboard. For entries marked as ’Context,’ only the highest AUROC is reported due to the absence of AUPRC metrics in the original publication. Abbreviations: CEHR-GPT-L (CEHR-GPT with linear probing), CEHR-GPT-T (CEHR-GPT with fine-tuning), GPT-L (The CEHR-GPT original variant with linear probing), GPT-T (The CEHR-GPT original variant with fine-tuning). Prev (Prevalence), R (AUROC), P (AUPRC)
@@ 5 Discussion
CEHR-GPT is, to our knowledge, the first EHR foundation model comprehensively evaluated across three core capabilities: feature representation, zero-shot prediction, and synthetic data generation. While prior models typically specialize in one area, CEHR-GPT demonstrates broad versatility, enabling a wide range of clinical tasks without requiring task-specific retraining. This flexibility is particularly valuable in real-world settings with limited labeled data, evolving populations, or rapidly changing outcome definitions.
A key enabler of this versatility is CEHR-GPT’s use of time tokens, along with Time Decomposition (TD) and Time-to-Event (TTE) objectives. These mechanisms help encode temporal structure directly into patient sequences and appear to enhance generalizability—particularly in zero-shot and transfer settings. For example, in the external ehrshot benchmark, CEHR-GPT achieved strong performance through vocabulary expansion and full fine-tuning. It matched top-reported models on Patient Outcome tasks and outperformed all previously reported models on New Diagnosis tasks. While MOTOR, CLMBR, Llama, and Mamba were evaluated using linear probing in prior studies, their performance could improve with fine-tuning.
In contrast to MOTOR, which consistently outperformed other models in linear probing, CEHR-GPT’s representation performance was more modest. This may reflect differences in pretraining: MOTOR was trained to forecast thousands of future events across time intervals, promoting long-range structure learning. CEHR-GPT, by comparison, was trained with next-token prediction and auxiliary temporal objectives—favoring sequential modeling over embedding quality. These results suggest that the alignment between pretraining and downstream tasks plays a critical role in determining the effectiveness of linear probing. Another factor is feature coverage: CEHR-GPT currently excludes measurements and observations, domains used by MOTOR. Expanding CEHR-GPT’s input domains may help close this gap.
Interestingly, GPT (CEHR-GPT without time-specific learning objectives) demonstrated similar performance to CEHR-GPT on in-distribution representation tasks, but underperformed in zero-shot prediction, treatment pathway analysis using synthetic data, and external validation. This suggests that time tokens act as a form of structural regularization—limiting overfitting to the training distribution while improving temporal generalization. Analogous to regularization in traditional models, the TD and TTE objectives guide the model to learn more transferable temporal representations, which is critical for generalization across health systems.
@@
Taken together, these results position CEHR-GPT as a flexible and general-purpose foundation model for EHR data. Its support for zero-shot prediction, synthetic data generation, and representation learning makes it well suited for diverse clinical workflows, including cohort discovery, surveillance, and rapid model prototyping. Future work may extend CEHR-GPT by incorporating additional domains (e.g., measurements, labs, observations), improving computational efficiency (e.g., via adapters or distillation), and enhancing clinical interpretability through human-in-the-loop evaluation.

### v2

@@ 4.4 Generalization on EHR-SHOT Benchmark
We demonstrated generalizability of CEHR-XGPT by evaluating it on the external ehrshot benchmarking dataset, which enables comparison with several competitive models available on the ehrshot Leaderboard, including MOTOR, Context, CLMBR, GBM, LR, RF, and GPT (Table 6 and Supplementary table 12). Overall, CEHR-XGPT outperformed GPT in both fine-tuned and linear-probing settings. For Patient Outcomes, fine-tuning CEHR-XGPT improved performance, with average gains of 1.2% in AUROC and 5.0% in AUPRC compared to linear probing. In contrast, on New Diagnosis tasks, fine-tuning CEHR-XGPT yielded a higher AUROC with an average gain of 1.9% but slightly reduced AUPRC by 0.6%. Interestingly, GPT followed a different pattern: for Patient Outcomes, fine-tuning and linear probing performed similarly, with fine-tuning achieving a modest average improvement of 0.9% in AUPRC. However, for New Diagnosis, fine-tuning degraded performance relative to linear probing, particularly for two low-prevalence cohorts, Celiac and Pancreatic Cancer. A similar trend was observed for CEHR-XGPT on Pancreatic Cancer, where linear probing outperformed fine-tuning. These results suggest that under extreme class imbalance, linear probing may offer a more robust approach for predicting rare events than fine-tuning.
More importantly, both fine-tuned and linear-probing versions of CEHR-XGPT demonstrated strong performance across these tasks, achieving the best results on New Diagnosis prediction compared to all other models. Notably, CEHR-XGPT outperformed models such as CLMBR, despite the latter being trained on data drawn from the same distribution (Stanford EHR data) as the ehrshot dataset. This suggests that simply extending the context window or adopting alternative architectures does not necessarily yield better performance on these tasks, and that CEHR-XGPT’s novel patient representation design and time-specific learning objectives may be key contributors to its improved performance.
@@
Table 6: Performance metrics of selected models for patient outcome and new diagnosis predictions using the ehrshot dataset. This table summarizes both individual and average model performance across tasks. The Best Baseline column shows the best metric among CLMBR, GBM, LR, and RF. For MOTOR, only aggregated performance data is available since specific task details were not published on the ehrshot leaderboard. For entries marked as ’Context,’ only the highest AUROC is reported due to the absence of AUPRC metrics in the original publication. Abbreviations: CEHR-XGPT-L (CEHR-XGPT with linear probing), CEHR-XGPT-T (CEHR-XGPT with fine-tuning), GPT-L (The CEHR-XGPT original variant with linear probing), GPT-T (The CEHR-XGPT original variant with fine-tuning). Prev (Prevalence), R (AUROC), P (AUPRC)
@@ 5 Discussion
CEHR-XGPT is, to our knowledge, the first EHR foundation model comprehensively evaluated across three core capabilities: feature representation, zero-shot prediction, and synthetic data generation. While prior models typically specialize in one area, CEHR-XGPT demonstrates broad versatility, enabling a wide range of clinical tasks without requiring task-specific retraining. This flexibility is particularly valuable in real-world settings with limited labeled data, evolving populations, or rapidly changing outcome definitions.
A key enabler of this versatility is CEHR-XGPT’s use of time tokens, along with Time Decomposition (TD) and Time-to-Event (TTE) objectives. These mechanisms help encode temporal structure directly into patient sequences and appear to enhance generalizability—particularly in zero-shot and transfer settings. For example, in the external ehrshot benchmark, CEHR-XGPT achieved strong performance through vocabulary expansion and full fine-tuning. It matched top-reported models on Patient Outcome tasks and outperformed all previously reported models on New Diagnosis tasks. While MOTOR, CLMBR, Llama, and Mamba were evaluated using linear probing in prior studies, their performance could improve with fine-tuning.
In contrast to MOTOR, which consistently outperformed other models in linear probing, CEHR-XGPT’s representation performance was more modest. This may reflect differences in pretraining: MOTOR was trained to forecast thousands of future events across time intervals, promoting long-range structure learning. CEHR-XGPT, by comparison, was trained with next-token prediction and auxiliary temporal objectives—favoring sequential modeling over embedding quality. These results suggest that the alignment between pretraining and downstream tasks plays a critical role in determining the effectiveness of linear probing. Another factor is feature coverage: CEHR-XGPT currently excludes measurements and observations, domains used by MOTOR. Expanding CEHR-XGPT’s input domains may help close this gap.
Interestingly, GPT (CEHR-XGPT without time-specific learning objectives) demonstrated similar performance to CEHR-XGPT on in-distribution representation tasks, but underperformed in zero-shot prediction, treatment pathway analysis using synthetic data, and external validation. This suggests that time tokens act as a form of structural regularization—limiting overfitting to the training distribution while improving temporal generalization. Analogous to regularization in traditional models, the TD and TTE objectives guide the model to learn more transferable temporal representations, which is critical for generalization across health systems.
@@
Taken together, these results position CEHR-XGPT as a flexible and general-purpose foundation model for EHR data. Its support for zero-shot prediction, synthetic data generation, and representation learning makes it well suited for diverse clinical workflows, including cohort discovery, surveillance, and rapid model prototyping. Future work may extend CEHR-XGPT by incorporating additional domains (e.g., measurements, labs, observations), improving computational efficiency (e.g., via adapters or distillation), and enhancing clinical interpretability through human-in-the-loop evaluation.